John deposited 3.2. 65690 € in a bank account which had an annual interest rate of 2 % (p.a.). The deposit was withdrawn 23.6. Calculate the net accrued interest when the tax at source rate was 30 %.



1
Expert's answer
2022-02-07T12:57:38-0500

If it was deposited on 3rd Feb and withdrawn on 23rd June on a non-reap year, with 365 days


Total days"=140"

Accrued interest"=\frac{2}{100}×\frac{1}{365}×140×65690"


"=503.9"


Net accrued interest "=\frac{70}{100}×503.9"

"=352.7\u20ac"